OpenStudy (anonymous):

Thermal physics! how do you rearrange the formula (q=mc ΔT) to the temp when you have everything except for the second temp, for the ΔT.

OpenStudy (anonymous):

\[q=mc(T_f-T_i)\] \[\frac{q}{mc}=T_f-T_i\] @Katelyn96 , which temperature do you have the final or the initial

OpenStudy (anonymous):

I only have the initial I think, its 20 degrees

OpenStudy (anonymous):

\[\frac{q}{mc}+T_i=T_f\]

OpenStudy (anonymous):

\[q=mcT_f-mcT_i\] \[q+mcT_i=mcT_f\] \[\frac{q}{mc}+T_i=T_f\]

OpenStudy (anonymous):

that is another way

OpenStudy (anonymous):

Okay, and this will work out the final temperature? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, thank you so much for your help.